. On January 2 of the current year for $300,000 Depreciator purchases new equipment for use in her business. The purchase is made from an unrelated person. The equipment has a 6-year class life and is 5-year property under § 168(c). Depreciator plans to use the equipment for seven years, and expects it to have a salvage value of $30,000 at the end of that time. Depreciator is a single, calendar year taxpayer, and she uses the equipment only in her business. In the following problems, compute the depreciation deductions with respect to the equipment in each year of its use and Depreciator’s adjusted basis for the property each year. (A) Depreciator does not elect out of Section 168K
Answer #1

usefull life of the equipment is 6years .so we have to depriciate the equipment for 6years.

under stright line depriciation method(SLM)

depriciation =(purchase price-salvage value)/use full life

=(300000-30000)/6

= 270000/6

depreciation for each year is =45000

adjested basis for the property each year are as follows

after completing year 1 value of equipment is = 300000-45000

=255000

year 2 =255000-45000

=210000

year 3 =210000-45000

=165000

year 4 =165000-45000

=120000

year 5 =120000-45000

=75000

year 6 =75000-45000

=30000

* 30000 is solvage value of the equipment.